OpenStudy (anonymous):

Given QRS=TUV, QS=3v+2 and TV=7v-6, find the length of QS and TV. A. 2 B. 9 C. 8 D. 20

OpenStudy (anonymous):

Have you drawn a picture yet?

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

try drawing one and see what you get

OpenStudy (anonymous):

Okay :)

OpenStudy (anonymous):

i'm assuming QRS and TUV are triangles. are they?

OpenStudy (anonymous):



OpenStudy (anonymous):

since the problem says that the triangles are equal we can say that 3v+2=7v-6 because of CPCTC

OpenStudy (anonymous):

And then what do i do?

OpenStudy (anonymous):

if we solve 3v+2=7v-6 we get v=2. If you plug 2 back into the original equations 3v+2 and 7v-6 we get they equal 8. So your answer is C 8. Did that make sense?

OpenStudy (anonymous):

@SkiTTleoooo47 did that make sense?

OpenStudy (anonymous):

Yes! Thank you :)

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

@habibmatatta did you get your other question asnswered? " what number exceeds its..."?

OpenStudy (anonymous):

Yeah i got it :) I understood :P